What volume will 12.0 g of oxygen gas (O,)

occupy at 25°C and a pressure of 52.7 kPa?


1
Expert's answer
2022-01-18T21:44:01-0500

Moles of oxygen gas ( O2 ) = n = 12 / 32 .

= 0.375 mol.

T = 25 degree celcius = 298 K .


P = 52.7 kpa = ( 52.7 × 1000 ) pa = 52700 pa .


PV = nR T .


V = nRT / P = (0.375 × 8.314 × 298) / (52700)


V = 0.0177 m3 = 17.7 litre